# Icy or Rocky? Convective or Stable?

### v1

@@ 7 Conclusions
  • Bridging empirical and physical interior modelling
@@
  • Uranus and Neptune as ”rock giants”
    We find that different bulk compositions are consistent with current data for both Uranus and Neptune. They encompass rock-dominated and water-rich interiors, with rock-to-water ratios varying by more than an order of magnitude (0.04–3.92 for Uranus, 0.20–1.78 for Neptune; Table 2). We hence confirm that the label ”ice giants” for Uranus and Neptune may be more of a historical artifact rather than a robust physical classification.
  • Differences between Uranus and Neptune
@@
  • Dynamo and immiscibility constraints matched
@@ Appendix A Additional interior data
Figure 8: The Brunt-Väisälä frequencies (top), entropies (middle), and temperature (bottom) profiles of the four Uranus models. For comparison, we also show purely adiabatic temperature profiles from Nettelmann et al. (2013) (dashed grey).
@@
We show interior temperature, entropy, and frequency profiles for the models of Uranus and Neptune in Figures 8 and 10. We find that, as expected, the Brunt-Väisälä frequency $N$ is nearly zero ($|N^{2}|<10^{-8}\text{h}^{-2}$) in adiabatic regions. This is orders of magnitudes smaller than analogous results in Morf et al. (2024). Note that $N^{2}$ should be smaller than zero throughout all adiabatic regions subjected to convection. This is violated in the outermost parts (normalised radii above 0.9) in Figures 8 and 10. However, these are caused by numerical floating point precision limitations with near-zero numbers, and they do not contain any physical meaning. The wiggles visible in Ledoux-stable regions ($N^{2}>10^{-2}\text{h}^{-2}$) originate from the random nature of the algorithm.
We show the relative pressure deviation criterion introduced in Equation 8 in Figure 9 for Uranus and Figure 11 for Neptune. We find that all the models are consistent with the $\epsilon=0.02$ limit introduced in Section 2. Note that in this work, the number $k$ in Equation 8 ranges from 3 to 8.
@@ Appendix B Compositional algorithm
We now discuss the compositional algorithm in detail, a part of the global algorithm depicted in Figure B.1. The compositional algorithm is depicted in Figure B.2 and represents an improved version of the work presented in Morf et al. (2024). As input, it requires both density values $\rho^{0},\dots,\rho^{N}]$ and pressure values $P^{0},\dots,P^{N}]$ within the planet. The index 0 denotes the outermost layer, where $N$ corresponds to the planetary centre. In this work, we define the first layer that has a pressure above 100 bars as the outermost layer. For layers below a pressure of 100 bar, we do not infer a composition and just fix the profile to follow the density-pressure atmosphere models proposed by Hueso et al. (2020). This is motivated by more complex chemical processes in the atmosphere (such as in Hueso et al. 2020), which our approach is not well suited for.
We consider hydrogen ($X$) and helium ($Y$), water ($Z_{1}$), rocks ($Z_{2}$), and iron ($Z_{3}$) for a composition mass fraction $\@vec{X}_{\text{EoS}}=(X,Y,Z_{1},Z_{2},Z_{3})$. To that end, we rely on the EoS provided by More et al. (1988); Chabrier & Debras (2021); Haldemann et al. (2020). The abundance of hydrogen and helium relative to each other is always fixed at a proto solar mass ratio of $0.705/0.275$. We assume the ideal mixing equation
@@
In the beginning, the compositional algorithm chooses a random temperature $T_{\text{EoS}}^{0}\in[250,450]$ K for the first layer that has a pressure above 100 bars. This temperature range is based on the results of Hueso et al. (2020). Given $T_{\text{EoS}}^{0}$, it tries to find a composition $\@vec{X}_{\text{EoS}}^{0}$ that results in a density $\rho_{\text{EoS}}^{0}$ that is close enough to the provided density $\rho^{0}$, meaning
@@
Here, we use $\delta=0.01$. If multiple such $\rho_{\text{EoS}}^{0}$ exist, we choose one of them randomly.
Moving towards the planetary interior, the compositional algorithm prioritizes finding regions of constant composition. These regions are adiabatic due to large scale convection mixing the materials homogeneously. To find them, the compositional algorithm checks whether it can keep the composition $\@vec{X}_{\text{EoS}}^{*}$ from the previous layer and follow an adiabatic gradient for the next $\Delta$ layers. The check is successful if the densities inferred by the EoS obey Equation 11 for all $\Delta$ layers. Note that we only accept such regions if $\Delta>N/8=64$. We introduce this limit to prevent small convection cells. While small convection cells are possible, sustaining them over long enough time scales is tricky and highly dependent on dynamical processes such as oscillatory double-diffusive convection (Leconte & Chabrier 2012; Tulekeyev et al. 2024). We do not investigate such dynamical processes better treated by evolution models (such as Tejada Arevalo 2025; Eberlein & Helled 2025), which motivates the aforementioned limit.
As a second priority, the compositional algorithm checks if it can find a homogeneous and adiabatic region by ignoring Equation 11 for $\Delta-1$ layers. Note that only satisfying Equation 11 again for the $\Delta$-th layer violates self-consistency to a high degree. However, due to the iterative nature of the global algorithm, the necessity for such violations should become less and less frequent.
Finally, as a third priority, the algorithm allows the composition to change: In addition to the inferred $\rho_{\text{EoS}}$ satisfying Equation 11, we now need to check for stability against convection. This is done by calculating the Brunt-Väisälä frequency $N$. For an arbitrary density $\rho$, pressure $P$, and temperature $T$, NN is given by (for example Unno et al. 1989; Kippenhahn et al. 2013)
@@
where the molecular weight $\mu$ parametrizes the composition. In particular, $N$ depends on the difference between the adiabatic gradient $\nabla_{\text{ad}}$ and temperature gradient $\nabla_{T}$
@@
taken in the limit $\delta\ln\rho\rightarrow 0$. $B$ can be approximated numerically (see Appendix B of Morf et al. 2024). The Brunt-Väisälä frequency $N$ denotes the oscillation frequency of a parcel of material that is slightly displaced compared to its surroundings while not exchanging any heat with it. For stability, we require $N^{2}\geq 0$. The criterion $N^{2}\geq 0$ is called the Ledoux criterion. If multiple tuples $\left(\rho^{i}_{\text{EoS}},T^{i}_{\text{EoS}},\@vec{X}^{i}_{\text{EoS}}\right)$ satisfy both Equation 11 and $N^{2}\geq 0$, we choose one of them randomly.
@@ Appendix C Algorithm convergence
Figure C.3: Maximal gravitational moment deviation vs. maximal relative pressure deviation for different generations $k\in\{1,3,5\}$ and in total over ten thousand Neptune models. Histograms of the depicted points are included for both axes. Points falling within the hatched area represent physical models. The arrows highlight one single path between generations.
@@
for different generations $k\in\{1,3,5\}$. Each data point represents one model, a complete tuple according to Equation 7. Points within the hatched area are both consistent with the $\epsilon=0.02$ criterion (see Equation 8) and with the measured gravitational moments within their respective uncertainties (see Table 1). Such points therefore correspond to physical models.
Figure C.3 highlights two properties of our algorithm: First, the histograms show that as we progress to higher generations, our models become more physical. They align more with the measured gravitational moments and also become more self-consistent according to Equation 8. Second, considering the single highlighted path (black arrows) between the generations, the random nature of our compositional algorithm becomes apparent. Note that while the models become more physical on average, that must not be true for any given model. Consider the first model (blue point with black circle, $k=1$) with pressure deviations of almost 100% and gravitational moments more than ten standard deviations away from the measured values. After just two generations (orange point with black circle, $k=3$), it leads to a physical model thanks to random choices made by the compositional algorithm. However, two additional generations later (green point with black circle, $k=5$), the result is again far away from the measured gravitational moments.
We can theretofore conclude that while the non-random parts of the algorithm lead towards physical models on average, the random nature of other algorithm parts ensure that a wide variety of options are explored. The aforementioned parameter $\epsilon$ should be thought of as a selection criterion, not a convergence criterion. There is no reason to prefer, for example, a generation $k=5$ models over a generation $k=3$ model. Higher generations are not necessarily refined versions of what came before. Finally, we note that Figure C.3 demonstrates that one could easily choose lower values for $\epsilon$. However, as discussed earlier, lower values for $\epsilon$ would not translate to a closer approach to physical reality. Each limitation listed in Section 6 is capable of introducing deviations comparable to or easily surpassing our $\epsilon=0.02$ threshold.
@@ Appendix D Comparing different EoS for SiO2
Throughout this work, we have emphasized that the EoS employed carry inherent uncertainties, which can lead to discrepancies when compared with alternative EoS formulations. To illustrate this, Figure D.4 presents a comparison focused solely on the SiO2 EoS, where we contrast the EoS used in this study (More et al. 1988) with the ANEOS formulation for forsterite (Stewart et al. 2019), constructed using M-ANEOS which is based on fitting analytic expressions of the Helmholtz free energy (Thompson et al. 1974; Melosh 2007; Thompson et al. 2019).
Figure D.4: Pressure vs. density of SiO2 for our Uranus and Neptune models. The solid lines correspond to the EoS of More et al. (1988), while the dashed lines to the ANEOS EoS. The density of SiO2 $(\rho_{Z_{2}})$ is not equal to the model density $(\rho)$ since there are no regions with pure rock in the models (see Equation 10). Note that we only show the regions where SiO2 is present in our models.
We find that the differences in density between the two SiO2 EoS are generally below 10%, with many models showing even smaller deviations. These differences only affect the SiO2 component in our ideal mixing formulation (Equation 10). When comparing pressures at constant density instead by inverting the relation, discrepancies in pressure can reach up to 50%. Our adopted consistency threshold of $\epsilon=0.02$ in Equation 8 is clearly conservative in comparison. Finally, we note that while both EoSs are developed for pure SiO2, the behaviour of materials in mixtures can differ significantly (immiscibility, phase transitions, etc.). It is therefore desirable to perform detailed simulations of different mixtures at planetary conditions. Currently, there is no clear justification for preferring one EoS over the other, especially for materials ”heavier” than hydrogen and helium.

### v2

@@ 7 Conclusions
  • Bridging empirical and physical interior modelling:
@@
  • Uranus and Neptune as ’rock giants’:
    We find that different bulk compositions are consistent with current data for both Uranus and Neptune. They encompass rock-dominated and water-rich interiors, with rock-to-water ratios varying by more than an order of magnitude (0.04–3.92 for Uranus, 0.20–1.78 for Neptune; Table 2). We hence confirm that the label ’ice giants’ for Uranus and Neptune may be more of a historical artifact rather than a robust physical classification.
  • Differences between Uranus and Neptune:
@@
  • Dynamo and immiscibility constraints matched:
@@ Appendix A Additional interior data
Figure 8: Brunt-Väisälä frequencies (top), entropies (middle), and temperature (bottom) profiles of the four Uranus models. For comparison, we also show purely adiabatic temperature profiles from Nettelmann et al. (2013) (dashed grey).
@@
We show interior temperature, entropy, and frequency profiles for the models of Uranus and Neptune in Figures 8 and 10. We find that, as expected, the Brunt-Väisälä frequency $N$ is nearly zero ($|N^{2}|<10^{-8}\text{h}^{-2}$) in adiabatic regions. This is orders of magnitudes smaller than analogous results in Morf et al. (2024). The frequency $N^{2}$ should be smaller than zero throughout all adiabatic regions subjected to convection. This is violated in the outermost parts (normalised radii above 0.9) in Figures 8 and 10. However, these are caused by numerical floating point precision limitations with near-zero numbers, and they do not contain any physical meaning. The wiggles visible in Ledoux-stable regions ($N^{2}>10^{-2}\text{h}^{-2}$) originate from the random nature of the algorithm.
We show the relative pressure deviation criterion introduced in Equation 8 in Figure 9 for Uranus and Figure 11 for Neptune. We find that all the models are consistent with the $\epsilon=0.02$ limit introduced in Section 2. In this work, the number $k$ in Equation 8 ranges from 3 to 8.
@@ Appendix B Compositional algorithm
We now discuss the compositional algorithm in detail, a part of the global algorithm depicted in Figure B.1. The compositional algorithm is depicted in Figure B.2 and represents an improved version of the work presented in Morf et al. (2024). As input, it requires both density values $\rho^{0},\dots,\rho^{N}]$ and pressure values $P^{0},\dots,P^{N}]$ within the planet. The index 0 denotes the outermost layer, where $N$ corresponds to the planetary centre. In this work, we defined the first layer that has a pressure above 100 bars as the outermost layer. For layers below a pressure of 100 bar, we did not infer a composition and just fixed the profile to follow the density-pressure atmosphere models proposed by Hueso et al. (2020). This is motivated by more complex chemical processes in the atmosphere (such as in Hueso et al. 2020), which our approach is not well suited for.
We considered hydrogen ($X$) and helium ($Y$), water ($Z_{1}$), rocks ($Z_{2}$), and iron ($Z_{3}$) for a composition mass fraction $\@vec{X}_{\text{EoS}}=(X,Y,Z_{1},Z_{2},Z_{3})$. To that end, we relied on the EoS provided by More et al. (1988); Chabrier & Debras (2021); Haldemann et al. (2020). The abundance of hydrogen and helium relative to each other was always fixed at a proto-solar mass ratio of $0.705/0.275$. We assumed the ideal mixing equation
@@
In the beginning, the compositional algorithm chooses a random temperature $T_{\text{EoS}}^{0}$. For this work, we randomly chose $T_{\text{EoS}}^{0}\in[250,450]$ K for the first layer that has a pressure above 100 bars. This temperature range is based on the results of Hueso et al. (2020). Given $T_{\text{EoS}}^{0}$, the compositional algorithm tries to find a composition $\@vec{X}_{\text{EoS}}^{0}$ that results in a density $\rho_{\text{EoS}}^{0}$ that is close enough to the provided density $\rho^{0}$, meaning
@@
Here, we used $\delta=0.01$. If multiple such $\rho_{\text{EoS}}^{0}$ exist, the algorithm chooses one of them randomly.
Moving towards the planetary interior, the compositional algorithm prioritises finding regions of constant composition. These regions are adiabatic due to large scale convection mixing the materials homogeneously. To find them, the compositional algorithm checks whether it can keep the composition $\@vec{X}_{\text{EoS}}^{*}$ from the previous layer and follow an adiabatic gradient for the next $\Delta$ layers. The check is successful if the densities inferred by the EoS obey Equation 11 for all $\Delta$ layers. We only accepted such regions if $\Delta>N/8=64$. This limit was introduced to prevent small convection cells. While small convection cells are possible, sustaining them over long enough time scales is tricky and highly dependent on dynamical processes such as oscillatory double-diffusive convection (Leconte & Chabrier 2012; Tulekeyev et al. 2024). We do not investigate such dynamical processes better treated by evolution models (such as Tejada Arevalo 2025; Eberlein & Helled 2025), which motivated the aforementioned limit.
As a second priority, the compositional algorithm checks if it can find a homogeneous and adiabatic region by ignoring Equation 11 for $\Delta-1$ layers. Only satisfying Equation 11 again for the $\Delta$-th layer violates self-consistency to a high degree. However, due to the iterative nature of the global algorithm, the necessity for such violations should become less and less frequent.
Finally, as a third priority, the algorithm allows the composition to change: In addition to the inferred $\rho_{\text{EoS}}$ satisfying Equation 11, one now needs to check for stability against convection. This is done by calculating the Brunt-Väisälä frequency $N$. For an arbitrary density $\rho$, pressure $P$, and temperature $T$, NN is given by (for example Unno et al. 1989; Kippenhahn et al. 2013)
@@
where the molecular weight $\mu$ parametrises the composition. In particular, $N$ depends on the difference between the adiabatic gradient $\nabla_{\text{ad}}$ and temperature gradient $\nabla_{T}$
@@
taken in the limit $\delta\ln\rho\rightarrow 0$. The mixing term $B$ can be approximated numerically (Appendix B of Morf et al. 2024). The Brunt-Väisälä frequency $N$ denotes the oscillation frequency of a parcel of material that is slightly displaced compared to its surroundings while not exchanging any heat with it. For stability, one requires $N^{2}\geq 0$. The criterion $N^{2}\geq 0$ is called the Ledoux criterion. If multiple tuples $\left(\rho^{i}_{\text{EoS}},T^{i}_{\text{EoS}},\@vec{X}^{i}_{\text{EoS}}\right)$ satisfy both Equation 11 and $N^{2}\geq 0$, the algorithm chooses one of them randomly.
@@ Appendix C Algorithm convergence
Figure C.3: Maximal gravitational moment deviation versus maximal relative pressure deviation for different generations $k\in\{1,3,5\}$ and in total over ten thousand Neptune models. Histograms of the depicted points are included for both axes. Points falling within the hatched area represent physical models. The arrows highlight one single path between generations.
@@
for different generations $k\in\{1,3,5\}$. Each data point represents one model, a complete tuple according to Equation 7. Points within the hatched area are both consistent with the $\epsilon=0.02$ criterion (Equation 8) and with the measured gravitational moments within their respective uncertainties (Table 1). Such points therefore correspond to physical models.
Figure C.3 highlights two properties of our algorithm: First, the histograms show that as we progress to higher generations, our models become more physical. They align more with the measured gravitational moments and also become more self-consistent according to Equation 8. Second, considering the single highlighted path (black arrows) between the generations, the random nature of our compositional algorithm becomes apparent. While the models become more physical on average, that must not be true for any given model. The first model (blue point with black circle, $k=1$) has pressure deviations of almost 100% and gravitational moments more than ten standard deviations away from the measured values. After just two generations (orange point with black circle, $k=3$), it leads to a physical model thanks to random choices made by the compositional algorithm. However, two additional generations later (green point with black circle, $k=5$), the result is again far away from the measured gravitational moments.
We can theretofore conclude that while the non-random parts of the algorithm lead towards physical models on average, the random nature of other algorithm parts ensure that a wide variety of options are explored. The aforementioned parameter $\epsilon$ should be thought of as a selection criterion, not a convergence criterion. There is no reason to prefer, for example, a generation $k=5$ models over a generation $k=3$ model. Higher generations are not necessarily refined versions of what came before. Finally, Figure C.3 demonstrates that one could easily choose lower values for $\epsilon$. However, as discussed earlier, lower values for $\epsilon$ would not translate to a closer approach to physical reality. Each limitation listed in Section 6 is capable of introducing deviations comparable to or easily surpassing our $\epsilon=0.02$ threshold.
@@ Appendix D Comparing different EoS for SiO2
Throughout this work, we have emphasised that the EoS employed carry inherent uncertainties, which can lead to discrepancies when compared with alternative EoS formulations. To illustrate this, Figure D.4 presents a comparison focused solely on the SiO2 EoS, where we contrast the EoS used in this study (More et al. 1988) with the ANEOS formulation for forsterite (Stewart et al. 2019), constructed using M-ANEOS which is based on fitting analytic expressions of the Helmholtz free energy (Thompson et al. 1974; Melosh 2007; Thompson et al. 2019).
Figure D.4: Pressure versus density of SiO2 for our Uranus and Neptune models. The solid lines correspond to the EoS of More et al. (1988), while the dashed lines to the ANEOS EoS. The density of SiO2 $(\rho_{Z_{2}})$ is not equal to the model density $(\rho)$ since there are no regions with pure rock in the models (Equation 10). We only show the regions where SiO2 is present in our models.
We find that the differences in density between the two SiO2 EoS are generally below 10%, with many models showing even smaller deviations. These differences only affect the SiO2 component in our ideal mixing formulation (Equation 10). When comparing pressures at constant density instead by inverting the relation, discrepancies in pressure can reach up to 50%. Our adopted consistency threshold of $\epsilon=0.02$ in Equation 8 is clearly conservative in comparison. Finally, while both EoSs are developed for pure SiO2, the behaviour of materials in mixtures can differ significantly (immiscibility, phase transitions, etc.). It is therefore desirable to perform detailed simulations of different mixtures at planetary conditions. Currently, there is no clear justification for preferring one EoS over the other, especially for materials ’heavier’ than hydrogen and helium.